(3 radical 5+ 2 radical 3)(3 radical 5 -2 radical 3)
Answer by edjones(8007)

Let sqrt(5)=f and sqrt(3)=t
(3f+2t)(3f-2t)
=9f^2-4t^2
=9*5-4*3
=45-12
=33
.
Ed